# Towards Understanding Valuable Preference Data for Large Language Model Alignment

### v1

@@ 5.2.1 Main Performance of Loss-IRM for Preference Data Selection
LossDiff-IRM achieves better performance. Table 3 summarizes the performance of Loss-IRM and competitors with DPO and SLiC. Across diverse LLM families, benchmarks and metrics, it can be observed that LossDiff-IRM occupies more darker cells, and surpasses full-data training while using only about 50%–65% of the data, indicating stronger performance. Specifically, compared to full-data training, LossDiff-IRM achieves average WinRate improvements of $13.50\%$, $9.76\%$, $16.17\%$, $18.02\%$ and $17.92\%$ on Llama-3.1-8B, Qwen3-8B-Base, and Pythia-2.8B/1.4B/410M with DPO, respectively. Furthermore, we compute exact TIF on smaller Pythia-410M. Table 4 shows that training on data selected by exact TIF and by LossDiff-IRM achieves comparable performance. These results demonstrate the effectiveness and superiority of Loss-IRM in selecting real valuable preference pairs that are really beneficial for current model alignment training.
Preference data quality is model-dependent. We observe that LossDiff-IRM outperforms selection based on Random, GPT4 score or external reward model, with only a few single-case exceptions, e.g., Qwen3-8B-Base on Vicuna-Bench under SLiC. Under DPO, relative to the second-best baseline, LossDiff-IRM delivers average WinRate improvements of $4.29\%$, $3.14\%$, $6.99\%$, $10.13\%$ and $6.40\%$ on Llama-3.1-8B, Qwen3-8B-Base, and Pythia-2.8B/1.4B/410M, respectively. Notably, GPT4 score selection can even reduce performance, such as on Qwen3-8B-Base with DPO. Moreover, Figure 4 shows that the Overlap Coefficient between selections varies across models; pairs within the same family (e.g., Pythia) exhibit higher overlap than cross-family pairs. Altogether, these findings indicate that preference data quality is inherently model dependent, and static or model-agnostic selectors may benefit one model yet harm another.
@@ A.1 Supplement to Preliminary
Reinforcement Learning with Human Feedback (RLHF). RLHF aims to align the given policy model $\pi_{\theta}$ with human preference by optimizing the model to maximize the expected reward value obtained from the reward model. The reward model is trained on a preference dataset, quantifying the human preference into scalar values. Typically, the Bradley-Terry (BT) model (Bradley & Terry, 1952) is used to estimate the probability distribution that a chosen response $y_{w}$ is preferred over a rejected one $y_{l}$ as follows:
@@
where $r(x,y)$ denotes the latent reward function, and $\sigma(\cdot)$ is the Sigmoid function. Due to unobservability of the reward function, the traditional RLHF typically follows a two-stage pipeline: a reward model $r_{\phi}(x,y)$ is pretrained on preference data; then, the policy model $\pi_{\theta}$ is updated based on the pretrained reward model using reinforcement learning (RL) algorithm like PPO Schulman et al. (2017). The overall objective of RLHF can be formulated as follows:
$\mathcal{L}_{\text{RLHF}}(\theta;\mathcal{D})=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]+\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)],$ (11)
@@ A.2 Related Work
LLM Preference Alignment. LLM preference alignment aims to steer LLM behaviors toward responses that better reflect human preferences and values (Wang et al., 2024). A common approach is RLHF that trains a reward model to provide reward signals for RL training (Bai et al., 2022b). Though effective, the two-stage RLHF pipeline is complex and resource-intensive, suffering from issues like reward hacking (Miao et al., 2024) and unstable optimization (Zheng et al., 2023b). To address these limitations, recent studies (Rafailov et al., 2023; Azar et al., 2024; Zhao et al., 2023; Meng et al., 2024; Ethayarajh et al., 2024) have proposed multiple alignment learning objectives that bypass explicit reward modeling, offering a simpler formulation for direct LLM preference alignment, such as DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), SLiC (Zhao et al., 2023; Liu et al., 2024), SimPO (Meng et al., 2024), KTO (Ethayarajh et al., 2024), CPO (Xu et al., 2024) and more (Wu et al., 2024b, a; Liu et al., 2024). Among them, DPO serves as a milestone work that derives the close-form expression of the optimal policy and substitutes it into the Bradley–Terry (BT) model, effectively hiding reward learning within the policy optimization process. SLiC adopts a hinge loss to enlarge the margin between the chosen and rejected responses. SimPO simplifies the DPO formulation and achieves reference model-free optimization. KTO employ prospect theory to directly maximizes the utility of generations rather than log-likelihood of preferences. Overall, while many efforts have been devoted to improving alignment algorithms, relatively little attention  (Deng et al., 2025; Gao et al., 2025) has been given to understanding the preference data quality, especially from model perspective.
@@
Data Selection for LLMs. Data selection (Albalak et al., 2024) aims to filter out low-quality or noisy data and retain high-quality data for better training and generalization, which plays a pivotal role in enhancing the performance and efficiency of LLMs across difference training stages, including pretraining (Penedo et al., 2023; Tang et al., 2024), supervised fine-tuning (SFT) (Pang et al., 2025; Qin et al., 2024), and preference alignment (Shen et al., 2024; Gao et al., 2025). Most existing approaches assess data from multiple considerations, such as importance (Xie et al., 2023) and diversity (Zhang et al., 2024a), using a variety of metrics such as text length (Nagatsuka et al., 2023), perplexity (Kong et al., 2024), text embeddings (Saranathan et al., 2024), or external signals from humans (Wu et al., 2023) or ChatGPT (Chen et al., 2024). We notice that most studies for LLM data selection primarily focus on the first two stages, i.e., pretraining and SFT, whereas the selection of preference data for alignment has received comparatively limited exploration. Existing studies (Pattnaik et al., 2024; Morimura et al., 2024; Deng et al., 2025) often employ off-the-shelf LLMs or pretrained reward models to pre-process preference data. For example, CurriDPO (Pattnaik et al., 2024) leverages the GPT-4 or reward model scores to organize curriculum learning. fDPO (Morimura et al., 2024) integrates a reward model into the DPO training process, filtering preference data based on reward scores. Muldrew et al. (2024) and Shen et al. (2025) incorporate active learning into alignment to improve data quality and annotation efficiency. Overall, these methods share a common trait: they rely on external signals–whether from powerful GPT models, pretrained reward models, or active human labeling–and thereby treat data quality as an inherent property of the data itself, while overlooking the role of the model and training objectives. Motivated by this limitation, we examine preference data quality from the model’s perspective to better understand which data are truly valuable for alignment.
@@ B.1 Formulations of SLiC Loss
Instantiation of SLiC Loss for Influence Function. By taking SLiC loss into the influence function formulation in Eq. (5), the SLiC-based influence function is derived as follows:
@@ B.2 Relationship Between Influence Influnction and LossDiff
Figure 7: Analysis of IF-based data partition with overlapping splits on Qwen3-0.6B-Base and Llama-3.2-1B. From Left to Right: Trainin DPO loss, evaluation DPO loss and evaluation reward margin. Subsets of {small, medium, large}-60% are denoted by {blue, orange, green}, respectively, while different line styles indicate IF values computed using different epoch checkpoints.
@@ D.2 Training Details
We conduct experiments using various LLM families, including Llama-3.1-8B (Vavekanand & Sam, 2024), Qwen3-8B-Base (Yang et al., 2025) and the Pythia series (Pythia-2.8B/1.4B/410M) (Biderman et al., 2023). Due to GPU limit, Llama-3.1-8B and Qwen3-8B-Base are trained using LoRA with $r_{\mathrm{LoRA}}=32,\alpha_{\mathrm{LoRA}}=32$ and $\mathrm{drop\_out=0.05}$, whereas the Pythia models are trained in a full-parameter setting. As described in the previous section, we first perform one epoch of SFT on each model based on UltraChat-200K dataset to initialize the alignment learning.
@@ D.3 Evaluation Details
Figure 11: Ablation study on Single Score. Comparison of LossDiff-IRM and its ablated variants that select data relying solely on LossDiff (“w/ LossDiff”) or solely on IRM (“w/ IRM”). Top: Training with DPO. Bottom: Training with SLiC.
@@
LLM DPO
Dataset
Ratio
UltraFeedback AlpacaEval Vicuna-Bench
@@
LLM SLiC
Dataset
Ratio
UltraFeedback AlpacaEval Vicuna-Bench
@@ E.2 More Visualization of IF
LLM
Training
Data
Method UltraFeedback AlpacaEval Vicuna-Bench
@@ E.3 Results of Direct Training on Train, Val, or Train+Val Sets
LLM
Validation
Dataset
DPO UltraFeedback AlpacaEval Vicuna-Bench

### v2

@@ 5.2.1 Main Performance of Loss-IRM for Preference Data Selection
LossDiff-IRM achieves better performance. Table 3 summarizes the performance of Loss-IRM and competitors with DPO and SLiC. Across diverse LLM families, benchmarks and metrics, it can be observed that LossDiff-IRM occupies more darker cells, and surpasses full-data training while using only about 50%–65% of the data, indicating stronger performance. Specifically, compared to full-data training, LossDiff-IRM achieves average WinRate improvements of $11.42\%$, $15.14\%$, $16.63\%$, $18.28\%$ and $17.71\%$ on Llama-3.1-8B, Qwen3-8B-Base, and Pythia-2.8B/1.4B/410M with DPO, respectively. Furthermore, we compute exact TIF on smaller Pythia-410M. Table 5 shows that training on data selected by exact TIF and by LossDiff-IRM achieves comparable performance. These results demonstrate the effectiveness and superiority of Loss-IRM in selecting real valuable preference pairs that are really beneficial for current model alignment training.
LossDiff-IRM outperforms several data-centric baselines. Table 4 reports the performance comparisons of LossDiff-IRM with several existing data-centric methods. It can be observed that our LossDiff-IRM outperforms these competitors on the most cases. Specifically, the average improvement of WinRate achieves $2.62\%$ over the best baselines across two models, especially achieving improvements of $6.45\%$ on average on Llama-3.8-8B. These results suggest the effectiveness of our IF-driven analysis and LossDiff-IRM data selection strategy. Additionally, compared to CurriDPO relied on external signals (GPT4 or reward model scores), $M_{\text{AP}}$ and RS-DPO, which adopt implicit reward margins or generate model-specific preference pairs using the SFT model and thus partially incorporate the model’s perspective, emerge as strong competitors. This aligns with the underlying philosophy of our idea: valuable preference data are model-dependent, rather than relying on external heuristics. The cost analysis between LossDiff-IRM and RS-DPO is provided in Appendix E.5.
Preference data quality is model-dependent. We observe that LossDiff-IRM outperforms selection based on Random, GPT4 score or external reward model, with only a few single-case exceptions, e.g., Qwen3-8B-Base on Vicuna-Bench under SLiC. Under DPO, relative to the second-best result, LossDiff-IRM delivers average WinRate improvements of $4.07\%$, $3.84\%$, $8.28\%$, $10.29\%$ and $8.13\%$ on Llama-3.1-8B, Qwen3-8B-Base, and Pythia-2.8B/1.4B/410M, respectively. Notably, GPT4 score selection can even reduce performance, such as on Qwen3-8B-Base with DPO. Moreover, Figure 5 shows that the Overlap Coefficient between selections varies across models; pairs within the same family (e.g., Pythia) exhibit higher overlap than cross-family pairs. Altogether, these findings indicate that preference data quality is inherently model dependent, and static or model-agnostic selectors may benefit one model yet harm another.
@@
5.2.2 Further Analysis
Figure 6: Analysis of percentile thresholds $\xi_{\text{small}},\xi_{\text{large}},\tau_{\text{small}},\tau_{\text{large}}$ of Llama-3.1-8B. We vary $\xi_{\text{small}}=(100-\xi_{\text{large}})$ and $\tau_{\text{small}}=(100-\tau_{\text{large}})$ with in $\{0,10,20,30\}$. Further analysis of Qwen3-8B-Base is illustrated in Figure 14 and Appendix E.6.
@@ 5.2.2 Further Analysis
Analysis of percentile thresholds $\xi_{\text{small}},\xi_{\text{large}},\tau_{\text{small}},\tau_{\text{large}}$. The percentile thresholds of LossDiff-IRM are tuned as hyperparameters. We vary $\xi_{\text{small}}=(100-\xi_{\text{large}})$ and $\tau_{\text{small}}=(100-\tau_{\text{large}})$ with in $\{0,10,20,30\}$. Figure 6 illustrates the performance curves varying with different thresholds for Llama-3.1-8B (DPO) across four benchmarks. It can be observed a rough trend: performance first improves and then degrades as the thresholds become stricter, that is, as more data are filtered out. This observation is expected, because initially, the threshold helps remove low-quality data; however, beyond a certain point, the filtering starts to exclude informative and high-quality data, leading to performance degradation. More analysis on Qwen3-8B-Base is provided in Appendix E.6.
@@ 6 Conclusion
Acknowledgments
ZZZ, QZW, JNZ and BH were supported by RGC General Research Fund No. 12200725, RIKEN Collaborative Research Fund, and HKBU CSD Departmental Incentive Scheme. MS was supported by JST ASPIRE Grant Number JPMJAP25B1. JCY is supported by National Natural Science Foundation of China (No. 62306178) and STCSM (No. 22DZ2229005).
Ethics Statement
This study adheres to the Code of Ethics. It relies only on publicly available models and datasets, without the use of sensitive or privacy data, and poses no identifiable risks concerning privacy, security, or fairness. The work is conducted purely for scientific purposes, and no conflicts of interest are involved.
Reproducibility Statement
We are committed to ensure the reproducibility of our proposed method. The detailed descriptions of our approach and experimental settings are all provided in this paper. The corresponding source code is released at https://github.com/tmlr-group/TIF_LossDiff-IRM. Both backbone models and datasets used in our work are publicly available. Furthermore, all parameters, hyper-parameters, and procedural steps required to reproduce our results are thoroughly recorded in the Experimental Details section and corresponding Appendix. We believe that these components provide the community with details necessary to verify and reproduce our work.
@@ A.1 Supplement to Preliminary
Reinforcement Learning with Human Feedback (RLHF). RLHF aims to align the given policy model $\pi_{\theta}$ with human preference by optimizing the model to maximize the expected reward value obtained from the reward model (Zhang et al., 2025). The reward model is trained on a preference dataset, quantifying the human preference into scalar values. Typically, the Bradley-Terry (BT) model (Bradley and Terry, 1952) is used to estimate the probability distribution that a chosen response $y_{w}$ is preferred over a rejected one $y_{l}$ as follows:
@@
where $r(x,y)$ denotes the latent reward function, and $\sigma(\cdot)$ is the Sigmoid function. Due to unobservability of the reward function, the traditional RLHF typically follows a two-stage pipeline: a reward model $r_{\phi}(x,y)$ is pretrained on preference data; then, the policy model $\pi_{\theta}$ is updated based on the pretrained reward model using reinforcement learning (RL) algorithm like PPO (Schulman et al., 2017). The overall objective of RLHF can be formulated as follows:
$\mathcal{L}_{\text{RLHF}}(\theta;\mathcal{D})=-\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)\right]+\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)],$ (11)
@@ A.2 Related Work
LLM Preference Alignment. LLM has achieved remarkable success across a wide range of tasks (Li et al., 2026; Chen et al., 2026). LLM preference alignment aims to steer LLM behaviors toward responses that better reflect human preferences and values (Wang et al., 2024). A common approach is RLHF that trains a reward model to provide reward signals for RL training (Bai et al., 2022b; Zhou et al., 2025a; b). Though effective, the two-stage RLHF pipeline is complex and resource-intensive, suffering from issues like reward hacking (Miao et al., 2024) and unstable optimization (Zheng et al., 2023b; Zhang et al., 2026). To address these limitations, recent studies (Rafailov et al., 2023; Azar et al., 2024; Zhao et al., 2023; Meng et al., 2024; Ethayarajh et al., 2024) have proposed multiple alignment learning objectives that bypass explicit reward modeling, offering a simpler formulation for direct LLM preference alignment, such as DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), SLiC (Zhao et al., 2023; Liu et al., 2024), SimPO (Meng et al., 2024), KTO (Ethayarajh et al., 2024), CPO (Xu et al., 2024) and more (Wu et al., 2024b; a; Liu et al., 2024). Among them, DPO serves as a milestone work that derives the close-form expression of the optimal policy and substitutes it into the Bradley–Terry (BT) model, effectively hiding reward learning within the policy optimization process. SLiC adopts a hinge loss to enlarge the margin between the chosen and rejected responses. SimPO simplifies the DPO formulation and achieves reference model-free optimization. KTO employ prospect theory to directly maximizes the utility of generations rather than log-likelihood of preferences. Overall, while many efforts have been devoted to improving alignment algorithms, relatively little attention  (Deng et al., 2025; Gao et al., 2025) has been given to understanding the preference data quality, especially from model perspective.
@@
Data Selection for LLMs. Data selection (Albalak et al., 2024) aims to filter out low-quality or noisy data (Zhou et al., 2024) and retain high-quality data for better training and generalization, which plays a pivotal role in enhancing the performance and efficiency of LLMs across difference training stages, including pretraining (Penedo et al., 2023; Tang et al., 2024), supervised fine-tuning (SFT) (Pang et al., 2025; Qin et al., 2024), and preference alignment (Shen et al., 2024; Gao et al., 2025). Most existing approaches assess data from multiple considerations, such as importance (Xie et al., 2023) and diversity (Zhang et al., 2024a), using a variety of metrics such as text length (Nagatsuka et al., 2023), perplexity (Kong et al., 2024), text embeddings (Saranathan et al., 2024), or external signals from humans (Wu et al., 2023) or ChatGPT (Chen et al., 2024). We notice that most studies for LLM data selection primarily focus on the first two stages, i.e., pretraining and SFT, whereas the selection of preference data for alignment has received comparatively limited exploration. Existing studies (Pattnaik et al., 2024; Morimura et al., 2024; Deng et al., 2025) often employ off-the-shelf LLMs or pretrained reward models to pre-process preference data. For example, CurriDPO (Pattnaik et al., 2024) leverages the GPT-4 or reward model scores to organize curriculum learning. fDPO (Morimura et al., 2024) integrates a reward model into the DPO training process, filtering preference data based on reward scores. Muldrew et al. (2024) and Shen et al. (2025) incorporate active learning into alignment to improve data quality and annotation efficiency. Overall, these methods share a common trait: they rely on external signals–whether from powerful GPT models, pretrained reward models, or active human labeling–and thereby treat data quality as an inherent property of the data itself, while overlooking the role of the model and training objectives. Motivated by this limitation, we examine preference data quality from the model’s perspective to better understand which data are truly valuable for alignment.
@@ B.1 Formulations of SLiC Loss
Instantiation of SLiC Loss for Influence Score. By taking SLiC loss into the influence function formulation in Eq. (5), the SLiC-based influence function is derived as follows:
@@ B.2 Relationship Between Influence Influnction and LossDiff
B.3 Formal Analysis of the Reliability of Medium-IF Data Selection
We demonstrate that the reliability of selecting high‑IF data decreases as the quality of the validation data decreases. We provide the derivation pipeline as follows:
Suppose the validation dataset is corrupted w.p. $\eta$. Then, the expected validation gradient $v$ consists of the clean ($c$) and noisy ($n$) parts, following $g_{v}=(1-\eta)g_{c}+\eta g_{n}$. For a training sample $i$ with the gradient $g_{i}$, we have its influence weight as
$\omega_{i}=-g_{v}^{\top}g_{i}=-(1-\eta)g_{c}^{\top}g_{i}-\eta g^{\top}_{n}g_{i},$ (21)
[…]
$\rho=\frac{\text{Cov}(\omega^{*}_{i},\omega_{i})}{\sqrt{\text{Var}(\omega^{*}_{i})\text{Var}(\omega_{i})}}=\sqrt{{\sigma^{2}_{w}}/{\sigma^{2}_{w}+\tau^{2}}}.$ (22)
As observed, when the corruption rate $\eta$ increases, the correlation $\rho$ between the true and observed scores decreases, leading to a lower probability $p(i\in\mathcal{T}_{\text{true}}(k)|i\in\mathcal{T}_{\text{obs}}(k))$. Hence, the top‑$k$ sampling strategy becomes less reliable under high noise levels. Similar derivations hold for bottom‑$k$ sampling, making middle‑$k$ sampling tend to be more reliable.
Figure 8: Analysis of IF-based data partition with overlapping splits on Qwen3-0.6B-Base and Llama-3.2-1B. From Left to Right: Trainin DPO loss, evaluation DPO loss and evaluation reward margin. Subsets of {small, medium, large}-60% are denoted by {blue, orange, green}, respectively, while different line styles indicate IF values computed using different epoch checkpoints. Previous analysis refers to Section 3.2 and Figure 1.
@@ D.1 Dataset Details
  • Area-Hard [^5] https://huggingface.co/datasets/lmarena-ai/arena-hard-auto (Li et al., 2024): This is a benchmark containing 500 challenging prompts curated by BenchBuilder, which utilizes LLM judges to estimate human preferences.
@@ D.2 Training Details
We conduct experiments using various LLM families, including Llama-3.1-8B (Vavekanand and Sam, 2024), Qwen3-8B-Base (Yang et al., 2025) and the Pythia series (Pythia-2.8B/1.4B/410M) (Biderman et al., 2023). Due to GPU limit, Llama-3.1-8B and Qwen3-8B-Base are trained using LoRA with $r_{\mathrm{LoRA}}=32,\alpha_{\mathrm{LoRA}}=32$ and $\mathrm{drop\_out=0.05}$, whereas the Pythia models are trained in a full-parameter setting. As described in the previous section, we first perform one epoch of SFT on each model based on UltraChat-200K dataset to initialize the alignment learning.
@@ D.3 Evaluation Details
We compare our LossDiff-IRM with several data-centric preference alignment methods, including:
  • CurriDPO (Pattnaik et al., 2024) is a curriculum learning-based method, which orders preference pairs to organize curriculum learning based on various creteria, such as GPT4 score (CurriDPO-GPT4) and reward model score (CurriDPO-Reward Model).
  • $M_{\text{AP}}$ (Huang et al., 2025) is a margin-based preference data selection metric, called alignment potential, which integrates both explicit and implicit reward margins to quantify preference data.
  • RS-DPO (Khaki et al., 2024) combines rejection sampling (RS) with DPO to generate high-quality preference data for DPO training. RS-DPO requires an external reward model to quantify preference pairs during rejection sampling.
Figure 12: Ablation study on Single Score. Comparison of LossDiff-IRM and its ablated variants that select data relying solely on LossDiff (“w/ LossDiff”) or solely on IRM (“w/ IRM”). Top: Training with DPO. Bottom: Training with SLiC. The ablation of WinRate is provided in Figure 3.
@@
LLM DPO Dataset Ratio UltraFeedback AlpacaEval Vicuna-Bench
@@
LLM SLiC Dataset Ratio UltraFeedback AlpacaEval Vicuna-Bench
@@ E.2 More Visualization of IF
LLM Training Data Method UltraFeedback AlpacaEval Vicuna-Bench
@@ E.3 Results of Direct Training on Train, Val, or Train+Val Sets
LLM Validation Dataset DPO UltraFeedback AlpacaEval Vicuna-Bench
@@ E.4 Impact of Introducing More Validation Data
Table 14: Time cost analysis between RS-DPO and LossDiff-IRM.
Trained Model # Prompts RS-DPO LossDiff-IRM
Time Throughput (prompts/sec) Time Throughput (prompts/sec)
Llama-3.1-8B 48,908 4 h 34 min 2.97 2 h 14 min 6.08
Qwen3-8B-Base 48,908 4 h 50 min 2.81 2h 43 min 5.03
E.5 Time Cost Analysis between LossDiff-IRM and RS-DPO
High-quality preference data selection and generation are two types of data-centric methods. For example, RS-DPO (Khaki et al., 2024) focuses on generating new high-quality preference pairs through rejection sampling using the SFT model, whereas our LossDiff-IRM focuses on analyzing existing annotated preference pairs to identify valuable preference pairs. Specifically, RS-DPO requires an external reward model to evaluate the responses generated by the SFT model. Given $N$ prompts and $K$ sampled responses per prompt, RS-DPO needs to score $N\times K$ responses with the reward model in order to construct contrastive preference pairs. For LossDiff-IRM, it requires scoring all preference pairs with both the current model and an auxiliary model. For the same scale of $N$ preference pairs, LossDiff-IRM requires $N$ forward passes with the current model and $N$ forward passes with the auxiliary model, resulting in a total of $2N$ forward passes, where the two models share the same architecture and differ only in parameters. To facilitate a direct comparison, Table 14 reports the data processing time cost of RS-DPO (including data generation and rejection sampling) using the OpenAssistant-1.4B reward model with K=8, as well as the data-selection time of LossDiff-IRM. All experiments were conducted on a single H100-80GB GPU.
The dominant cost of RS-DPO comes from using the reward model to evaluate all $N\times K$ responses generated by the SFT model. Even using a relatively small 1.4B reward model, the time cost is actually higher than LossDiff-IRM. Therefore, the time cost of LossDiff-IRM is at least on par with RS-DPO, and even often smaller, especially when considering that RS-DPO scales with $K$ or uses a larger reward model.
Figure 14: Analysis of percentile thresholds $\xi_{\text{small}},\xi_{\text{large}},\tau_{\text{small}},\tau_{\text{large}}$ of Qwen3-8B-Base. We vary $\xi_{\text{small}}=(100-\xi_{\text{large}})$ and $\tau_{\text{small}}=(100-\tau_{\text{large}})$ with in $\{0,10,20,30\}$. Analysis for Llama-3.1-8B is illustrated in Figure 6.
Figure 15: Visualization of distribution of LossDiff and implicit reward margin for Llama-3.1-8B and Qwen3-8B-Base, respectively.
E.6 More Analysis of Percentile Thresholds
As a supplement, Figure 14 provides the analysis of percentile thresholds $\xi_{\text{small}},\xi_{\text{large}},\tau_{\text{small}},\tau_{\text{large}}$ for Qwen3-8B-Base. The rough trends are similar as for Llama-3.1-8B: performance first improves and then degrades as the thresholds become stricter. Following this observation, we finally set $\xi_{\text{small}}$ and $\xi_{\text{small}}$ as 10, $\tau_{\text{small}}$ and $\tau_{\text{small}}$ as 10 for all models.
E.7 Visualization of Distribution of LossDiff and IRM
Figure 15 visualizes the LossDiff and IRM distributions for Llama-3.1-8B and Qwen3-8B-Base, respectively. We observe that both LossDiff and IRM exhibit distributional shapes resembling unimodal, Gaussina-like distributions, where the majority of data lies in the middle region and only a small number of data fall in the extremes. In our data selection strategy, both criteria concentrate data in their middle ranges, taking their intersection would not result in an extremely small subset. The overlap between their medium regions remains sufficiently large and high-quality for alignment.